a loaded tube is 1.5kg placed in liquid of 9.5g/dm³ was made to oscillate vertically when slightly pushed down, if the time for 50 oscillation is 35 sec determine. a) periodic frequency b) angular velocity


1
Expert's answer
2021-03-08T08:24:25-0500

Mass of tube"(M)=1.5\space kg"

Density of tube "(\rho)=9.5\space g/dm^3=9.5 \space kg/m
^3"


(a) Time taken for 50 oscillations = 35 s

Time Period "(T)=\dfrac{35}{50}=0.7\space s"

Frequency"(f)=\dfrac{1}{T}=1.428\space s^{-1}"


(b) Angular Velocity"(\omega)=2\pi T=4.39\space rad/s"